# On Sobolev and Besov spaces with hybrid regularity

### v1

@@ 1. Introduction
This classical result holds if the space $H^{q}(\square)$ is discretised by isotropic wavelets. The recent articles [3, 19, 25] show, however, that the classical Sobolev spaces $H^{q}(\square)$ can also be characterised by tensor-product or hyperbolic wavelets. This approach leads to a new perspective on the best $N$-term approximation, as it is well-known that tensor-product wavelets can approximate functions essentially better. This concept is known as the Sparse Grid [2, 13]. Here, the term essentially is understood as $n$-times as well up to logarithmic terms.
@@
In [22, 23], Besov spaces have been used for the approximation with tensor-product wavelets in $L^{2}(\square)$ and $H^{1}(\square)$. Therein, the respective approximation spaces have been defined as as the tensor product of quasi-Banach spaces, resulting in function spaces of Besov type with hybrid regularity. If $q\geq 0$, this procedure can easily be extended to $H^{q}(\square)$, but whenever $q$ is strictly positive, the approximation spaces are of hybrid regularity. However, if one wishes to approximate a function in $H^{q}(\square)$ for negative $q$, which is necessary in the case of e.g. boundary integral equations, cf. [24, 28], these results do not carry over, as the resulting approximation spaces can no longer be written as the intersection of tensor-product spaces.
@@ 2. Wavelet Bases
In this section, we define the wavelet bases under consideration and state their most important properties. Throughout the article, we assume that the scaling functions and the wavelets involved are compactly supported with support sizes depending on their levels. In particular, we assume that $\operatorname{diam}(\operatorname{supp}\phi_{\lambda})\sim\operatorname{diam}(\operatorname{supp}\psi_{\lambda})\sim 2^{-|\lambda|}$ holds for any one-dimensional scaling function $\phi_{\lambda}$ and wavelet $\psi_{\lambda}$ on level $|\lambda|$. Moreover, we require that this property also holds for the dual scaling functions and the dual wavelets. The wavelets we are going to consider are, in general, biorthogonal and were first constructed in [4]. Later, this construction was also transferred to a finite interval in [7]. Alternatively, as shown in [5], it is also possible to use Daubechies wavelets [9, 10] on the interval. However, as for numerical applications there is a need for an efficient evaluation of the primal wavelets, we especially emphasise the constructions of [4, 7].
@@ 2.1. Univariate Wavelet Bases
[…]
@@
which are used to discretise a vector space $V$ consisting of functions (or distributions) on the unit interval $[0,1]$. Typically, $V=H^{q}([0,1])$ is a Sobolev space with regularity $q$. We assume that the function spaces $V_{j}$ can be generated by shifts and dyadic dilations of a scaling function $\phi$, with possible modifications at the endpoints of the interval. Moreover, for a suitable index set $\Delta_{j}$, we assume that
@@
[…]
@@ 2.1.1. Multiscale Bases on $[0,1]$
with the complement or difference space $W_{j}$. One can show that, if the scaling function $\phi$ generates a shift-invariant space, cf. [4, 7], there exist mother wavelet functions $\psi$ such that
@@
is a basis set of $W_{j}$. Also herein, $\nabla_{j}$ is a suitable index set and $\psi_{\lambda}$ is a properly scaled and translated copy of a mother wavelet (again with possible modifications at the endpoints of the interval). For convenience, if $\lambda\in\Delta_{j}$ or $\lambda\in\nabla_{j}$, let us denote its level by $|\lambda|\mathrel{\mathrel{\mathop{:}}=}j$. Furthermore, due to (1), both $\phi$ and all mother wavelets $\psi$ can be obtained by translated copies of the refined function $\phi$.
@@
provided that $j_{0}<j$. If we also define $W_{j_{0}}\mathrel{\mathrel{\mathop{:}}=}V_{j_{0}}$, $\nabla_{j_{0}}\mathrel{\mathrel{\mathop{:}}=}\Delta_{j_{0}}$, and $\psi_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\phi_{\lambda}$ for $\lambda\in j_{0}$, as the function spaces $V_{j}$ are asymptotically dense, we conclude that the set
@@
If we follow the construction of [4, 7], and use the appropriate scaling $\|\psi_{\lambda}\|_{L^{2}([0,1])}\sim 1$, the set $\Psi$ forms a Riesz basis of $L^{2}([0,1])$, meaning that
@@
The spaces $\tilde{V}_{j}\mathrel{\mathrel{\mathop{:}}=}\{\tilde{\phi}_{\lambda}:\lambda\in\Delta_{j}\}$ admit the regularity $\tilde{\gamma}>0$ and the approximation order $\tilde{d}$. This fact provides us the number of vanishing moments of the wavelet $\psi$, which means that $\langle p,\,\psi\rangle=0$ for any polynomial $p$ up to the order $\tilde{d}$. Finally, there exists also a unique biorthogonal wavelet basis $\tilde{\Psi}=\{\tilde{\psi}_{\lambda}:\lambda\in\nabla\}$, satisfying
@@
For the setting of piecewise constant wavelets with three vanishing moments $(d,\tilde{d})=(1,3)$ and piecewise linear wavelets with two vanishing moments $(d,\tilde{d})=(2,2)$, an illustration is provided in Figure 1 and Figure 2, respectively. Notice that in the latter figure, the primal wavelets depicted discretise functions with with zero boundary conditions at $x=0$ and nonzero boundary conditions at $x=1$.
@@ 2.1.2. Multiscale Transforms
Herein, the matrices $\mathbf{M}_{j,0},\tilde{\mathbf{M}}_{j,0}\in\mathbb{R}^{|\Delta_{j}|\times|\Delta_{j-1}|}$ and $\mathbf{M}_{j,1},\tilde{\mathbf{M}}_{j,1}\in\mathbb{R}^{|\Delta_{j}|\times|\nabla_{j}|}$ are called the refinement masks. By the local supports and the chosen scaling, the matrices $\big{[}\mathbf{M}_{j,0},\mathbf{M}_{j,1}\big{]}$ and $\big{[}\tilde{\mathbf{M}}_{j,0},\tilde{\mathbf{M}}_{j,1}\big{]}$ are uniformly stable.
@@
$u=\sum_{\lambda\in\Delta_{j}}c_{\lambda}\phi_{\lambda}=\sum_{\lambda\in\Delta_{j-1}}c_{\lambda}\phi_{\lambda}+\sum_{\lambda\in\nabla_{j}}d_{\lambda}\psi_{\lambda}.$
In view of (4), this is equivalent to
@@
where $\mathbf{c}_{j}\mathrel{\mathrel{\mathop{:}}=}[c_{\lambda}]_{\lambda\in\Delta_{j}}$, $\mathbf{d}_{j}\mathrel{\mathrel{\mathop{:}}=}[d_{\lambda}]_{\lambda\in\nabla_{j}}$. On the other hand, remarking that $\mathbf{c}_{j}=\langle\tilde{\Phi}_{j},u\rangle$ and $\mathbf{d}_{j}=\langle\tilde{\Psi}_{j},u\rangle$, we obtain that
@@
Hence, we conclude that $[\tilde{\mathbf{M}}_{j,0},\tilde{\mathbf{M}}_{j,1}]^{\intercal}[\mathbf{M}_{j,0},\mathbf{M}_{j,1}]=\mathbf{I}$, and therefore
@@
and accordingly, on the dual side,
@@ 2.2.1. Isotropic Wavelet Bases
Let us first address the bivariate case. It is trivial to deduce that for any $j\geq j_{0}$, the basis set $\Phi_{j}\otimes\Phi_{j}$ discretises the space $V_{j}\otimes V_{j}$. As for $j\geq j_{0}+1$ we also have
$V_{j}\otimes V_{j}=\big{(}V_{j-1}\otimes V_{j-1}\big{)}\oplus\big{(}V_{j-1}\otimes W_{j}\big{)}\oplus\big{(}W_{j}\otimes V_{j-1}\big{)}\oplus\big{(}W_{j}\otimes W_{j}\big{)},$
we can define
@@ 2.2.2. Tensor-Product Wavelet Bases
Another approach is to use the tensor product of one-dimensional wavelets on all the different levels. This approach is straightforward: for a given multiindex $\mathbf{j}\geq\mathbf{j}_{0}+\mathbf{1}$, we define the index set $\boldsymbol{\nabla}_{\mathbf{j}}\mathrel{\mathrel{\mathop{:}}=}\nabla_{j_{1}}\times\dots\times\nabla_{j_{n}}$, and for $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\boldsymbol{\nabla}_{\mathbf{j}}$, we define the wavelet function
@@
2.3. Auxillary Results
In this section, we state and prove three lemmata. When comparing the different function spaces in Section 5.2, they will be crucial.
@@ 2.3. Auxillary Results
$\|\mathbf{A}\|_{p}\leq\big{\|}\mathbf{A}^{\odot p}\big{\|}_{1}^{\frac{1}{p}}=\left[\max_{1\leq j\leq n}\sum_{i=1}^{m}\big{|}a_{i,j}\big{|}^{p}\right]^{\frac{1}{p}},$ (8)
@@
The subadditivity (9) implies that the function $\|\cdot\|_{p}^{p}$ is subadditive. Hence, the quasi-triangle inequality
$\|\mathbf{u}+\mathbf{v}\|_{p}\leq 2^{\frac{1}{p}}\big{(}\|\mathbf{u}\|_{p}+\|\mathbf{v}\|_{p}\big{)}$
follows, cf. [12].
@@
Let $m\geq j_{0}$ and consider the wavelet transforms $\mathbf{T}_{m}$ and $\tilde{\mathbf{T}}_{m}$ given by (6) and (7), respectively. Then, for $0<p\leq 2$, there holds
@@
We will show the theorem on the primal side only, as for the dual side, the same arguments can be used. First, we remark that, for $p=2$, both statements are true since both $\Phi$ and $\Psi$ are uniformly stable bases of $L^{2}([0,1])$.
@@
we fix $\mu\in\Delta_{m}$. Since $\phi_{\mu}$ is compactly supported, on each level $j_{0}\leq j\leq m$, there are at most $\mathcal{O}(1)$ wavelets $\psi_{\lambda}$ with $\operatorname{supp}\psi_{\lambda}\cap\operatorname{supp}\phi_{\mu}\neq\emptyset$, by which
$\|\mathbf{T}_{m}\|_{\infty}\lesssim\max_{\mu\in\Delta_{m}}\sum_{j=j_{0}}^{m}\max_{\lambda\in\nabla_{j}}|t_{\mu,\lambda}|.$
Since $\|\phi_{\mu}\|_{L^{\infty}}\sim 2^{\frac{m}{2}}\sim 2^{\frac{m-j}{2}}\|\psi_{\lambda}\|_{L^{\infty}}$, we conclude that for any $\lambda\in\nabla_{j}$, there must hold
$|t_{\mu,\lambda}|\lesssim 2^{\frac{j-m}{2}},$ (11)
by which
$\|\mathbf{T}_{m}\|_{\infty}\lesssim\sum_{j=j_{0}}^{m}2^{\frac{j-m}{2}}\lesssim 1.$
Let us finally treat the case $10<p<$ 1. By Lemma 2.1 and the arguments above, we have
$\|\mathbf{T}_{m}^{\intercal}\|_{p}^{p}\leq\max_{\mu\in\Delta_{m}}\sum_{j=j_{0}}^{m-1}\sum_{\lambda\in\nabla_{j}}\big{|}t_{\mu,\lambda}\big{|}^{p}\lesssim\sum_{j=j_{0}}^{m}2^{p\frac{j-m}{2}}\lesssim 1,$
by which $\|\mathbf{T}_{m}^{\intercal}\|_{p}$ is bounded independently from $m$ as well. This proves the first inequality.
To show the second inequality, we will first show that $\|\mathbf{T}_{m}\|_{1}\lesssim 2^{\frac{m}{2}}$. If this holds true, then for $\frac{1}{p}=\frac{1-\theta}{1}+\frac{\theta}{2}=1-\frac{\theta}{2}$, i.e., $11-\theta=\frac{2}{p}-$ 1, we can conclude by interpolation, that
@@
Let $j_{0}\leq j\leq m$ and $\lambda\in\nabla_{j}$. As the wavelet $\psi_{\lambda}$ satisfies $\operatorname{diam}(\operatorname{supp}\psi_{\lambda})\lesssim 2^{-j}$, and $\operatorname{diam}(\operatorname{supp}\phi_{\mu})\lesssim 2^{-m}$, there are at most $\mathcal{O}(2^{m-j})$ nontrivial coefficients $t_{\mu,\lambda}$ in (10). Hence, by (11), there holds
$\sum_{\mu\in\Delta_{m}}|t_{\mu,\lambda}|\lesssim 2^{m-j}\cdot 2^{\frac{j-m}{2}}=2^{\frac{m-j}{2}}\leq 2^{\frac{m}{2}},$
by which also $\|\mathbf{T}_{m}\|_{1}\lesssim 2^{\frac{m}{2}}$.
Finally, for $10<p<$ 1, we again make use Lemma 2.1, which implies that
$\|\mathbf{T}_{m}\|_{p}^{p}$ $\leq\max_{j_{0}\leq j\leq m}\max_{\lambda\in\nabla_{j}}\sum_{\mu\in\Delta_{m}}|t_{\mu,\lambda}|^{p}\lesssim\max_{j_{0}\leq j\leq m}2^{m-j}2^{p\frac{j-m}{2}}=2^{m(1-\frac{p}{2})}.$
After taking the $p$-th root, we arrive at $\|\mathbf{T}_{m}\|_{p}\lesssim 2^{m(\frac{1}{p}-\frac{1}{2})}$, which is what we wanted to show. ∎
Remark 2.4.
In the proof of Lemma 2.3, the crucial step is exploiting the compact support of the wavelets and their duals. As already stated, such wavelets exist and were constructed on the real line and adapted to the interval, cf. [5, 9, 10] for orthonormal wavelets and [4, 7] for the biorthogonal spline wavelets, which are numerically more feasible.
On the other hand, the criterion of the compact support may be relaxed: Indeed, what we reqire is that the row- and column sums (if $p<1$ for the $p$-th power) of the matrix can be controlled sufficiently well.

### v2

# On Sobolev and Besov spaces of hybrid regularity

@@ 1. Introduction
This classical result holds if the space $H^{q}(\square)$ is discretised by isotropic wavelets. The recent articles [3, 22, 33] show, however, that the classical Sobolev spaces $H^{q}(\square)$ can also be characterised by tensor-product or hyperbolic wavelets. This approach leads to a new perspective on the best $N$-term approximation, as it is well-known that tensor-product wavelets can approximate functions essentially better. This concept is known as sparse grid, cf. [2, 15]. Here, the term essentially is understood as $n$-times as well up to logarithmic terms.
@@
In [29, 30], Besov spaces have been used for the approximation with tensor-product wavelets in $L^{2}(\square)$ and $H^{1}(\square)$. Therein, the respective approximation spaces have been defined as the tensor product of quasi-Banach spaces, resulting in function spaces of Besov type of hybrid regularity. If $q\geq 0$, this procedure can easily be extended to $H^{q}(\square)$, but whenever $q$ is strictly positive, the approximation spaces are of hybrid regularity. However, if one wishes to approximate a function in $H^{q}(\square)$ for negative $q$, which is necessary in the case of e.g. boundary integral equations, cf. [31, 35], these results do not carry over, as the resulting approximation spaces can no longer be written as the intersection of tensor-product spaces.
@@ 2. Wavelet Bases
In this section, we define the wavelet bases under consideration and state their most important properties. Throughout the article, we assume that the scaling functions and the wavelets involved are compactly supported or decay sufficiently fast. Moreover, we require that this property also holds for the dual scaling functions and the dual wavelets. Numerically feasible wavelets are, in general, biorthogonal and were first constructed in [5]. These wavelets yield a compact support and therefore, $\operatorname{diam}(\operatorname{supp}\phi_{\lambda})\sim\operatorname{diam}(\operatorname{supp}\psi_{\lambda})\sim 2^{-|\lambda|}$ holds for any one-dimensional scaling function $\phi_{\lambda}$ and wavelet $\psi_{\lambda}$ on level $|\lambda|$. Later, this construction was also transferred to a finite interval in [8]. Alternatively, as shown in [6, 27], it is also possible to use Daubechies wavelets [10, 11] on the interval. However, as for numerical applications there is a need for an efficient evaluation of the primal wavelets, we especially emphasise the constructions of [5, 8].
@@ 2.1. Univariate Wavelet Bases
We consider a sequence of nested, finite-dimensional, and asymptotically dense function spaces
@@
which are used to discretise a vector space $V$ consisting of functions (or distributions) on the unit interval $I\mathrel{\mathrel{\mathop{:}}=}[0,1]$. Typically, $V=H^{q}([0,1])$ is a Sobolev space with regularity $q$. We assume that the function spaces $V_{j}$ can be generated by shifts and dyadic dilations of scaling functions, with possible modifications at the endpoints of the interval. Moreover, for a suitable index set $\Delta_{j}$, we assume that
@@
is a Riesz basis of $V_{j}$. Here, the index $\lambda=(j,k)$ contains information of the level $j$ and the location $k$. Let us mention that the spaces $(V_{j})_{j\geq j_{0}}$ form a multiresolution analysis on the unit interval $[0,1]$, see e.g. [25]. In the easiest case, one can think of $\phi$ as the constant function $1$, and of $\phi_{\lambda}$ as a properly scaled, dyadic indicator function, i.e., $\phi_{\lambda}=2^{\nicefrac{{j}}{{2}}}\mathds{1}_{[2^{-j}k,\,2^{-j}(k+1)]}=2^{\nicefrac{{j}}{{2}}}\mathds{1}_{I}(2^{j}x-k)$.
@@ 2.1.1. Multiscale Bases on $[0,1]$
with the complement or difference space $W_{j}$. One can show that, if the scaling function $\phi$ generates a shift-invariant space, cf. [5, 8], there exist wavelet functions such that
@@
is a basis set of $W_{j}$. Also herein, $\nabla_{j}$ is a suitable index set and $\psi_{\lambda}$ is a properly scaled and translated copy of a mother wavelet (again with possible modifications at the endpoints of the interval). For convenience, if $\lambda\in\Delta_{j}$ or $\lambda\in\nabla_{j}$, let us denote its level by $|\lambda|\mathrel{\mathrel{\mathop{:}}=}j$.
@@
provided that $j_{0}<j$. If we also define $W_{j_{0}}\mathrel{\mathrel{\mathop{:}}=}V_{j_{0}}$, $\nabla_{j_{0}}\mathrel{\mathrel{\mathop{:}}=}\Delta_{j_{0}}$, and $\psi_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\phi_{\lambda}$ for $\lambda\in\nabla_{j_{0}}$, as the function spaces $V_{j}$ are asymptotically dense, we conclude that the set
@@
Following [6, 8, 27] for example, the set $\Psi$ forms a Riesz basis of $L^{2}([0,1])$, meaning that
@@
The spaces $\tilde{V}_{j}\mathrel{\mathrel{\mathop{:}}=}\{\tilde{\phi}_{\lambda}:\lambda\in\Delta_{j}\}$ admit the regularity $\tilde{\gamma}>0$ and the approximation order $\tilde{d}$. This fact provides us the number of vanishing moments of the wavelets $\psi_{\lambda}$, which means that $\langle p,\,\psi_{\lambda}\rangle=0$ for any polynomial $p$ up to the order $\tilde{d}$. Finally, there exists also a unique biorthogonal wavelet basis $\tilde{\Psi}=\{\tilde{\psi}_{\lambda}:\lambda\in\nabla\}$, satisfying
@@
For the setting of [8, 9], i.e., compactly supported, piecewise constant wavelets with three vanishing moments $(d,\tilde{d})=(1,3)$ and piecewise linear wavelets with two vanishing moments $(d,\tilde{d})=(2,2)$, an illustration is provided in Figure 1 and Figure 2, respectively. Notice that in the latter figure, the primal wavelets depicted discretise functions with zero boundary conditions at $x=0$ and nonzero boundary conditions at $x=1$.
@@ 2.1.2. Multiscale Transforms
Herein, the matrices $\mathbf{M}_{j,0},\tilde{\mathbf{M}}_{j,0}\in\mathbb{R}^{|\Delta_{j}|\times|\Delta_{j-1}|}$ and $\mathbf{M}_{j,1},\tilde{\mathbf{M}}_{j,1}\in\mathbb{R}^{|\Delta_{j}|\times|\nabla_{j}|}$ are called the refinement masks. By the local supports and the chosen scaling, the matrices $\big[\mathbf{M}_{j,0},\mathbf{M}_{j,1}\big]$ and $\big[\tilde{\mathbf{M}}_{j,0},\tilde{\mathbf{M}}_{j,1}\big]$ are uniformly stable in the sense that their condition numbers are uniformly bounded.
@@
$u=\sum_{\lambda\in\Delta_{j}}c_{\lambda}\phi_{\lambda}=\sum_{\lambda\in\Delta_{j-1}}c_{\lambda}\phi_{\lambda}+\sum_{\lambda\in\nabla_{j}}d_{\lambda}\psi_{\lambda}.$ (6)
By interpreting the coefficients $c_{\lambda}$ and $d_{\lambda}$ as finite-dimensional column vectors, we may use (4) to deduce that (6) can equivalently be written as
@@
where $\mathbf{c}_{j}\mathrel{\mathrel{\mathop{:}}=}[c_{\lambda}]_{\lambda\in\Delta_{j}}$ and $\mathbf{d}_{j}\mathrel{\mathrel{\mathop{:}}=}[d_{\lambda}]_{\lambda\in\nabla_{j}}$. On the other hand, remarking that $\mathbf{c}_{j}=\langle\tilde{\Phi}_{j},u\rangle$ and $\mathbf{d}_{j}=\langle\tilde{\Psi}_{j},u\rangle$, we obtain that
@@
Hence, we conclude that $\tilde{\mathbf{M}}_{j,0},\tilde{\mathbf{M}}_{j,1}]^{\intercal}[\mathbf{M}_{j,0},\mathbf{M}_{j,1}]=\mathbf{I}$, and therefore
@@
and accordingly on the dual side
@@ 2.2.1. Isotropic Wavelet Bases
Let us first address the bivariate case. It is trivial to deduce that the basis set $\Phi_{j}\otimes\Phi_{j}$ discretises the space $V_{j}\otimes V_{j}$ for any $j\geq j_{0}$. As we also have
$V_{j}\otimes V_{j}=\big(V_{j-1}\otimes V_{j-1}\big)\oplus\big(V_{j-1}\otimes W_{j}\big)\oplus\big(W_{j}\otimes V_{j-1}\big)\oplus\big(W_{j}\otimes W_{j}\big)$
for $j\geq j_{0}+1$, we can define
@@ 2.2.2. Tensor-Product Wavelet Bases
Another approach is to use the tensor product of one-dimensional wavelets on all the different levels. This approach is straightforward: for a given multiindex $\mathbf{j}$ which satisfies the component-wise inequality $\mathbf{j}\geq\mathbf{j}_{0}+\mathbf{1}$, we define the index set $\boldsymbol{\nabla}_{\mathbf{j}}\mathrel{\mathrel{\mathop{:}}=}\nabla_{j_{1}}\times\dots\times\nabla_{j_{n}}$, and for $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{n})\in\boldsymbol{\nabla}_{\mathbf{j}}$, we define the wavelet function
@@
2.3. Auxiliary Results
In this section, we state and prove four lemmata. When comparing the different function spaces in Section 5.2, they will be crucial.
@@ 2.3. Auxiliary Results
$\|\mathbf{A}\|_{p}^{p}\leq\big\|\mathbf{A}^{\odot p}\big\|_{1}=\max_{1\leq j\leq n}\sum_{i=1}^{m}\big|a_{i,j}\big|^{p},$ (9)
@@
The subadditivity (10) implies that the function $\|\cdot\|_{p}^{p}$ is subadditive for $0<p\leq 1$, cf. [13]. As the function $x\mapsto x^{\nicefrac{{1}}{{p}}}$ is convex, the quasi-triangle inequality
$\|\mathbf{u}+\mathbf{v}\|_{p}\leq\big(\|\mathbf{u}\|_{p}^{p}+\|\mathbf{v}\|_{p}^{p}\big)^{\frac{1}{p}}=2^{\frac{1}{p}}\bigg(\frac{1}{2}\|\mathbf{u}\|_{p}^{p}+\frac{1}{2}\|\mathbf{v}\|_{p}^{p}\bigg)^{\frac{1}{p}}\leq 2^{\frac{1}{p}-1}\big(\|\mathbf{u}\|_{p}+\|\mathbf{v}\|_{p}\big)$
follows for all $\mathbf{u},\mathbf{v}\in\ell^{p}$.
To derive estimates on the wavelet transform matrices, we need to bound the matrix coefficients as well. The decay behaviour of the coefficients is governed by the decay behaviour of the wavelet itself.
@@
Suppose that the scaling functions and the wavelets satisfy the decay estimate
$|\phi_{\lambda}(x)|,|\psi_{\lambda}(x)|\lesssim 2^{\nicefrac{{j}}{{2}}}(1+|2^{j}x-k|)^{-\alpha},\qquad\lambda=(j,k),$ (11)
and likewise for the dual side for some $\alpha>1$. Then, for $j_{0}\leq j\leq m$ and any two indices $\lambda=(j,k)\in\nabla_{j}$ and $\mu=(m,\ell)\in\Delta_{m}$, there holds
$|t_{\mu,\lambda}|,\big|\tilde{t}_{\mu,\lambda}\big|\lesssim 2^{\frac{j-m}{2}}\big(1+|k-2^{j-m}\ell|\big)^{-\alpha}.$ (12)
Proof.
First, we note that
$t_{\mu,\lambda}$ $=\int_{I}\psi_{\lambda}(x)\tilde{\phi}_{\mu}(x)\operatorname{d}\!x.$
Since $\|\psi_{\lambda}\|_{L^{\infty}(I)}\lesssim 2^{\nicefrac{{j}}{{2}}}$ and $\big\|\tilde{\phi}_{\mu}\big\|_{L^{1}(I)}\lesssim 2^{-\nicefrac{{m}}{{2}}}$ by (11), this implies (12) for all pairs of $k$ and $\ell$ such that $|k-2^{j-m}\ell|\leq 1$.
If $|k-2^{j-m}\ell|\geq 1$, we distinguish between the following two cases. First, if $|2^{m}x-\ell|\geq 2^{\frac{m-j}{\alpha}-1}|k-2^{j-m}\ell|$, there follows
$|t_{\mu,\lambda}|$ $\lesssim 2^{\frac{j+m}{2}}\int_{I}\big|\psi_{\lambda}(x)\big|\big(1+|2^{m}x-\ell|\big)^{-\alpha}\operatorname{d}\!x$
$\lesssim 2^{\frac{m-j}{2}}\big(1+2^{\frac{m-j}{\alpha}-1}|k-2^{j-m}\ell|\big)^{-\alpha}$
$\lesssim 2^{\frac{j-m}{2}}|k-2^{j-m}\ell|^{-\alpha}.$
On the other hand, if $|2^{m}x-\ell|\leq 2^{\frac{m-j}{\alpha}-1}|k-2^{j-m}\ell|$, then we have
$|2^{j}x-k|$ $=\big|2^{j-m}(2^{m}x-\ell)+2^{j-m}\ell-k\big|\geq|2^{j-m}\ell-k|-2^{j-m}|2^{m}x-\ell|$
$\geq|2^{j-m}\ell-k|-2^{j-m}2^{\frac{m-j}{\alpha}-1}|k-2^{j-m}\ell|$
$=|2^{j-m}\ell-k|\bigg(1-\frac{1}{2}\cdot 2^{(m-j)(\frac{1}{\alpha}-1)}\bigg)$
$\geq\frac{1}{2}\cdot|2^{j-m}\ell-k|$
because $\alpha>1$. Therefore, we can conclude that
$|t_{\mu,\lambda}|$ $\lesssim 2^{\frac{j+m}{2}}\int_{I}\big(1+|2^{j}x-k|\big)^{-\alpha}\big|\tilde{\phi}_{\mu}(x)\big|\operatorname{d}\!x$
$\lesssim 2^{\frac{j-m}{2}}\big(1+|2^{j-m}\ell-k|\big)^{-\alpha}$
$\lesssim 2^{\frac{j-m}{2}}|k-2^{j-m}\ell|^{-\alpha}$
also in this case. The estimates on $\tilde{t}_{\mu,\lambda}$ follow in the same manner. ∎
Remark 2.4.
Lemma 2.3 relies on the decay condition (11) only. In particular, (11) holds whenever
$\phi_{\lambda}(x)=2^{\nicefrac{{j}}{{2}}}\phi(2^{j}x-k),\qquad|\phi(x)|\lesssim(1+|x|)^{-\alpha},$
which is true as soon as $\phi$ gives rise to a regular multiresolution analysis in the sense of [25]. Obviously, (11) holds also for wavelets with compact support on both the primal and dual side. Such wavelets have been constructed on the real line and adapted to the interval, cf. [6, 10, 11, 27] for orthonormal wavelets and [5, 8] for the biorthogonal spline wavelets. Additionally, there are smooth wavelets on the real line which decay faster than any polynomial, see [26].
Lemma 2.5.
Suppose that (11) holds for some $\alpha>1$ on the primal and dual side. Let $m\geq j_{0}$ and consider the wavelet transforms $\mathbf{T}_{m}$ and $\tilde{\mathbf{T}}_{m}$ given by (7) and (8), respectively. Then, for $\nicefrac{{1}}{{\alpha}}<p\leq 2$, there holds
@@
We will show this lemma on the primal side only, as the same arguments can be used for the dual side. First, we remark that, for $p=2$, both statements are true since both $\Phi$ and $\Psi$ are uniformly stable bases of $L^{2}([0,1])$.
@@
we fix $\mu=(m,\ell)\in\Delta_{m}$. In view of Lemma 2.3, there are at most $\mathcal{O}(1)$ wavelets $\psi_{\lambda}$, $\lambda=(j,k)\in\nabla_{j}$ with $|k-2^{j-m}\ell|\leq 1$, which implies that
$\sum_{j=j_{0}}^{m}\sum_{\lambda\in\nabla_{j}}|t_{\mu,\lambda}|$ $\lesssim\sum_{j=j_{0}}^{m}2^{\frac{j-m}{2}}\left[1+\sum_{\begin{subarray}{c}k\in\mathbb{Z}\\ |k-2^{j-m}\ell|\geq 1\end{subarray}}|k-2^{j-m}\ell|^{-\alpha}\right]$
$\lesssim\sum_{j=j_{0}}^{m}2^{\frac{j-m}{2}}$
$\lesssim 1$
as $\alpha>1$. Since $\mu\in\Delta_{m}$ was arbitrary, this implies the first estimate for all $1\leq p\leq 2$ by interpolation. If $p<1$, we use the same arguments to conclude that
$\sum_{j=j_{0}}^{m}\sum_{\lambda\in\nabla_{j}}|t_{\mu,\lambda}|^{p}\lesssim\sum_{j=j_{0}}^{m}2^{p\frac{j-m}{2}}\lesssim 1$
as $\alpha p>1$, so Lemma 2.1 yields the first estimate for $\nicefrac{{1}}{{\alpha}}<p\leq 1$.
To show the second inequality, we will first show that $\|\mathbf{T}_{m}\|_{1}\lesssim 2^{\nicefrac{{m}}{{2}}}$. If this holds true, then we can conclude for $\frac{1}{p}=\frac{1-\theta}{1}+\frac{\theta}{2}=1-\frac{\theta}{2}$, i.e., $1-\theta=\frac{2}{p}-1$, by interpolation that
@@
Let $j_{0}\leq j\leq m$ and $\lambda=(j,k)\in\nabla_{j}$. Again in view of Lemma 2.3, there are at most $\mathcal{O}(2^{m-j})$ indices $\mu=(m,\ell)\in\Delta_{m}$ satisfying $|k-2^{j-m}\ell|\leq 1$. Hence, there holds
$\sum_{\mu\in\Delta_{m}}|t_{\mu,\lambda}|$ $\lesssim 2^{\frac{j-m}{2}}\left[2^{m-j}+\sum_{\begin{subarray}{c}\ell\in\mathbb{Z}\\ |k-2^{j-m}\ell|\geq 1\end{subarray}}|k-2^{j-m}\ell|^{-\alpha}\right]$
$=2^{\frac{m-j}{2}}+2^{\frac{j-m}{2}}\sum_{\begin{subarray}{c}\ell\in\mathbb{Z}\\ |\ell|\geq 2^{m-j}\end{subarray}}2^{\alpha(m-j)}|\ell|^{-\alpha}$
$\lesssim 2^{\frac{m-j}{2}}+2^{(m-j)(\alpha-\frac{1}{2})}2^{(m-j)(1-\alpha)}$
$\sim 2^{\frac{m-j}{2}},$
by which $\|\mathbf{T}_{m}\|_{1}\lesssim 2^{\frac{m}{2}}$.
Finally, for $0<p<1$, we use similar arguments to obtain that
$\sum_{\mu\in\Delta_{m}}|t_{\mu,\lambda}|^{p}\lesssim 2^{(m-j)(1-\frac{p}{2})},\qquad\alpha p>1,$
which allows us to conclude the second estimate using Lemma 2.1. ∎